OpenStudy (anonymous):

The function f (x) = x^4 has an axis of symmetry the vertical axis y. True or False? This is true right??

OpenStudy (wolfe8):

I forgot how to check this, but I think you need to find where the gradient is = 0.

OpenStudy (anonymous):

One way to check is of course to plot the given function. You'll see it has an axis of symmetry at y-axis. Another way is to replace x with -x. If the resulting function is the same as the original, it has an axis of symmetry at the y-axis. Like, in the given function, f(x) = x^4 f(-x) = (-x)^4 f(-x) = x^4 = f(x) i think so...

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

do you agree?

OpenStudy (jdoe0001):

yes, you're correct on the check for f(-x) and you're also correct on checking the graph too, mind you that for any function of that kind with a EVEN exponent, y-axis is always their line of symmetry

OpenStudy (anonymous):

what about this one? In the values interval [0, 2p] the function f (x) = sin(x) has more periods by the function g (x) = sin(3x). True or False?

OpenStudy (anonymous):

i thought that was true but...i am not sure... now i believethat is false!!

OpenStudy (jdoe0001):

\(\bf sin(Bx) \textit{ period is } \cfrac{\textit{regular period of sine}}{B} \implies \cfrac{2\pi}{B}\\ sin(3x) \textit{ period is }\cfrac{2\pi}{3}\)

OpenStudy (jdoe0001):

so, sin(x) regular period is \(\bf 2\pi\) sin(3x) has \(\bf \cfrac{2\pi}{3}\) so 2/3 of a pi, one period 4/3 of a pi, 2 periods 6/3 of a pi, 3 periods notice that 6/3 = 2, that is \(\bf 2\pi\) so sin(3x) has 3 times the periods as sin(x)

OpenStudy (anonymous):

so you say that is true?

OpenStudy (jdoe0001):

well, notice, sin(3x) has 3 periods between 0 and \(2\pi\) sin(x) only has one, same interval

OpenStudy (anonymous):

yes so its true...

OpenStudy (anonymous):

oh...is the opposite...what i am thinking??? ahaha./.. it is false!!

OpenStudy (jdoe0001):

well, it says "has more periods by the function g(x)" it doesn't say "has more periods THAN the function g(x)"

OpenStudy (jdoe0001):

does sin(x) has more periods than sin(3x)? no does sin(x) has more periods if it were rewritten as sin(3x)? yes

OpenStudy (anonymous):

oh pellet... one word ... ....by...... means than the first has fewer from second?

OpenStudy (anonymous):

i am not fluent in english

OpenStudy (jdoe0001):

well, I guess we dunno what they mean... can you post a screenshot of it?

OpenStudy (anonymous):

hahaha it is not in english,,, ket me see my translation again it was my mistake..

OpenStudy (anonymous):

i used gogle translator... oh...

OpenStudy (anonymous):

the correct translation is : f(x) has more periods THAN the function g(x)

OpenStudy (anonymous):

so is False

OpenStudy (jdoe0001):

ohh, then is false :)

OpenStudy (anonymous):

just tell me... can i use ....by... when?

OpenStudy (anonymous):

for example 4 is bigger by 3?? can i say this?

OpenStudy (anonymous):

it asks if 3 is bigger than 4 but stating"is 4 bigger by 3??"

OpenStudy (anonymous):

btw they are both correct thank you!

OpenStudy (jdoe0001):

"by" denotes a relationship of dependency something by something else, means something is affected by something else "than" denotes comparison something than something else, means something in comparison to something else

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

yes i know for "than"

OpenStudy (anonymous):

i did not know about by

OpenStudy (anonymous):

so i have to join this site whenever i have time to improve my english ahahaha

OpenStudy (jdoe0001):

hehehe